vaporeon246:

For questions 1 and 2 use f(x) = x 5 and g(x) = 2x 1) Find (f plus g)(x) Question 1 answer choices: a) 3x 5 b) x 5 c) 2x 10 d) 2x^2 5 2) Find ( f x g ) (x) question 2 answer choices f) 2x^2 5 g) 3x^2 10x h) 2x^2 10x j) 2x 10

ThisGirlPretty:

@sillybilly123

sillybilly123:

\( f(x) = x^5\) and \(g(x) = 2x\) ??

sillybilly123:

scan the qu. if you can

vaporeon246:

It's algebra 2

dude:

1) Find (f g)(x) You mean \(f~օ~ g(x)\)?

dude:

Ah, and \(\color{#0cbb34}{\text{Originally Posted by}}\) @sillybilly123 \( f(x) = x^5\) and \(g(x) = 2x\) ?? \(\color{#0cbb34}{\text{End of Quote}}\)

vaporeon246:

but that isn't one of the answer choices

dude:

No, are those equations the ones in the question?

vaporeon246:

no f(x) is x plus 5 and g(x) is 2x

vaporeon246:

For some reason the plus sign keeps disappearing from the first equation

dude:

\[x+5+2x\] Simply combine like terms \(\color{red}x+5+\color{red}{2x}\)

vaporeon246:

so wouldn't it be 2x^2 + 5?

dude:

You're just adding, you would do that if you were multiplying :P

vaporeon246:

oh yeah lol

vaporeon246:

and #2 is 2x^ +5 right?

dude:

\(2x(x+5)\) Distribute \(\boxed{2x^2+10x}\)

vaporeon246:

Ok Thank you @dude

dude:

You're welcome :P

vaporeon246:

How do i put a photo on here?

sillybilly123:

screengrab etc on a lappie dunno about an app, doubt it exists. so email yourself a pic and give a URL and link it ...... in a browser....